40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6798     Entry time: Tue Jun 12 01:58:33 2012
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: aligned Y arm to Y end green 

[Jenne, Yuta]

We aligned Y arm to the Y end green incident beam.
We noticed two TEM00, bright and dim, so we decreased Y end laser temperature to 34.13 deg C.
It doubled the transmission of the green, and now the transmission to the PSL table is 178 uW, which is close to the maximum(197 uW) we got so far.

Current settings for Y end laser is;

  Y end laser "T+": 34.049 deg C
  Y end laser "ADJ": 0
  Y end laser measured temperature: 34.13 deg C
  C1:GCY-SLOW_SERVO2_OFFSET = 31025
  Y end slow servo: on (was off)

We aligned IR beam to the Y arm by mostly adjusting PZTs and got the transmission, C1:LSC-TRY_OUT ~ 0.9.

We tried to calculate the mode-matching ratio for IR by taking TRY data while ITMY and ETMY are swinging (without ALS), but it was difficult because we see too many higher order modes.

Tomorrow, we will (1) connect the beatbox to ADC, (2) edit c1gcv model, (3) scan the arm using I-Q signals.

ELOG V3.1.3-